VEGAS Team Meeting

February 06, 2012

Attendees

  • CASPER Team:
    • Dan Werthheimer
    • Mark Wagner

  • NRAO:
    • John Ford
    • Jason Ray
    • Randy McCullough
    • Mark Whitehead
    • Patrick Brandt
    • Galen Watts
    • Marty Bloss
    • Anish Roshi

Meeting Notes:

  • Logistics for March tests and observations.
    • Berkeley/Caltech could plan to come to Green Bank the WEEK OF MARCH 5 to work with NRAO on testing and integration, before the galactic center observations. Probably 3-4 people
  • March Testing Status
    • Mode 1 working and tested at high clock rate
      • Up to 357 MHz clock rate (2.85 GSamples/sec - 1.42 GHz bandwidth) with switching signals and logic. Good packets and spectrum. Changes will be documented on wiki this week. Binaries can be downloaded. The next incremental increase easier on R2 and mode 13 next
      • filter 1400 for mode #1
    • Computers for test
      • looking into using 8 computers for tests - if NICs can be found machines may be available between NRAO and UC-B - report next week.
    • IF/LO hardware
      • waiting for IF PCBs - in process, expected this week
    • ADC testing
      • Anish will do some tests w/synth this week and send around plots
      • Reviewed the checklist that Jason built for the hardware items for the March test.
      • There is a strong preference for the overall configuration to mimic the final configuration. Plan for eight Valons for the test at this point.
    • NRAO Modification Request (MR) status: Noted as 'VEGAS' in NRAO Plan of Record
    • Software integration tasks
      • Integrated testing of VEGAS manager with other GBT software and checking the data with GBTIDL.
        • Blocked by a firmware bug - access to new binaries will help
      • Testing the VEGAS Manager along with GBTIDL to see whether the state integration is done correctly.
        • Need a FITS file, new CASPER code, switching details, this weeks version should wotk
      • Testing VEGAS configuration for mode 1 (and 13) using config tool.
        • with CCC fo approval, Changes can be scheduled upon approval.
      • Data writing speed is waiting for a new host (coming with UC-B) before test is executed *Other status:
    • 15 Roach 2 prototypes are built. 11 work and they are looking at the broken ones. The problems are very distinct between the 4 that don't work so the problems are presumed to be from manufacture, not design. The plan is to update design to eliminate the blue wire changes and move forward with next run.
    • The mezzanine board would be useful but is not required for VEGAS. It is not working yet, but expected to be working soon.
    • Jason created a list transceivers, switches, and fibers. Orders sgould be placed for three (3) Netgear switches and three (3) tranceivers. The rest of the order can wait until the final Roach 2 and ADC order.

MartyBloss - 2012-02-06
Topic revision: r2 - 2012-02-06, MartyBloss
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding NRAO Public Wiki? Send feedback